OpenStudy (anonymous):

What is the Equivalent fraction of 8 over 15?

OpenStudy (ash2326):

We have \[\frac{8}{15}\] To find an equivalent fraction of any fraction either we divide numerator and denominator by a common factor. Or we could multiply numerator and denominator by a number. Here the fraction is in its lowest form. So we'll have to use the second method, let's multiply numerator and denominator by 2( can choose any number) \[\frac{8\times 2}{15\times 2}=\frac{16}{30}\] You can multiply by any number you want